OpenStudy (anonymous):

I need help finding all the solutions for sec x=√2

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ sec(\theta)=\cfrac{1}{cos(\theta)}}}\\ \quad \\ \quad \\ sec(x)=\sqrt{2}\implies \cfrac{1}{cos(x)}=\sqrt{2}\implies \cfrac{1}{\sqrt{2}}=cos(x)\\ \quad \\ {\color{blue}{ \cfrac{1}{\sqrt{2}}\cdot \cfrac{\sqrt{2}}{\sqrt{2}}\implies \cfrac{\sqrt{2}}{2} }}\qquad thus\\ \quad \\ \cfrac{1}{\sqrt{2}}=cos(x)\implies \cfrac{\sqrt{2}}{2}=cos(x)\) so... that's a well known angle, and should be in your Unit Circle

OpenStudy (anonymous):

thank you so much!

OpenStudy (jdoe0001):

yw